Solve for x
2(x-4)-(x+2)=(x+2)(x-4)

Answer by checkley77(12844)

2(x-4)-(x+2)=(x+2)(x-4)
2x-8-x-2=x^2-2x-8
x-10=x^2-2x-8
x^2-2x-8-x+10=0
x^2-3x+2=0
(x-2)(x-1)=0
x-2=0
x=2 answer
x-1=0
x=1 answer.